# MediConfusion: Can you trust your AI radiologist? Probing the reliability of multimodal medical foundation models

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have tremendous potential to improve the accuracy, availability, and cost-effectiveness of healthcare by providing automated solutions or serving as aids to medical professionals. Despite promising first steps in developing medical MLLMs in the past few years, their capabilities and limitations are not well-understood. Recently, many benchmark datasets have been proposed that test the general medical knowledge of such models across a variety of medical areas. However, the systematic failure modes and vulnerabilities of such models are severely underexplored with most medical benchmarks failing to expose the shortcomings of existing models in this safety-critical domain. In this paper, we introduce MediConfusion, a challenging medical Visual Question Answering (VQA) benchmark dataset, that probes the failure modes of medical MLLMs from a vision perspective. We reveal that state-of-the-art models are easily confused by image pairs that are otherwise visually dissimilar and clearly distinct for medical experts. Strikingly, all available models (open-source or proprietary) achieve performance below random guessing on MediConfusion, raising serious concerns about the reliability of existing medical MLLMs for healthcare deployment. We also extract common patterns of model failure that may help the design of a new generation of more trustworthy and reliable MLLMs in healthcare. The evaluation code and the dataset are available at our github.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated unprecedented capabilities in a variety of multimodal tasks, including image understanding and visual reasoning, autonomous driving (Cui et al., 2024), robotics (Wang et al., 2024a) and embodied AI (Driess et al., 2023). Motivated by this success, a growing body of work (Moor et al., 2023; Li et al., 2024; Lin et al., 2023) explores the potential of MLLMs in medical applications with the hope of paving the way to more accurate, personalized and cost-effective healthcare solutions through modern generative AI.
@@
Even though MLLMs show enormous potential in a wide range of tasks, a swath of challenges have stymied their deployment, including object hallucinations Li et al. , relationship hallucinations (Wu et al., 2024), inaccurate object counting (Jain et al., 2024) and lack of spatial reasoning capabilities (Kamath et al., 2023). These shortcomings are especially worrisome in safety-critical applications, such as healthcare, where reliability is an essential requirement. In fact, recent research efforts on medical MLLMs have revealed weak anatomic knowledge (Nan et al., 2024), concerns on toxicity and patient privacy Xia et al. (2024), highly unreliable disease diagnosis Wu et al. (2023a), and the fact that even a junior doctor far outperforms the most proficient medical MLLMs across a wide spectrum of tasks Wang et al. (2024b). As model failure in the medical domain can lead to serious adverse health effects, it is of utmost importance to understand the performance and limitations of generative AI in the medical context.
@@
A flurry of activity has emerged around probing the performance of medical MLLMs in a multitude of tasks, including visual question answering (VQA) (Ben Abacha et al., 2021), disease classification and report generation (Royer et al., 2024), and modality recognition (Wu et al., 2023b). Even though the proposed medical benchmarks offer valuable insights on model performance across a variety of anatomic regions and imaging modalities, they are focused on evaluating the medical knowledge of MLLMs across large evaluation sets, heavily biased towards common or typical scenarios. Therefore, it is unclear how well the measured performance correlates with the actual multimodal medical reasoning capabilities of these models, especially in the face of systematic but perhaps more intricate model failures underrepresented in the dataset. Therefore, developing new evaluation benchmarks that carefully test and probe the capabilities of these systems, expose their vulnerabilities, and facilitate the development of a better understanding of failure modes is vital in healthcare applications.
@@
In this work, we introduce MediConfusion, a challenging benchmark for evaluating the failure modes of medical MLLMs from a vision perspective. We combine novel insights on the image representations of medical MLLMs with the expertise of clinical radiologists to craft a benchmark dataset that stress-tests the visual capabilities of state-of-the-art models. In particular, our work reveals that medical MLLMs often confuse image pairs that otherwise appear very different in the image domain. Leveraging this observation, we introduce an automated pipeline to discover such pairs in the ROCO (Pelka et al., 2018) multimodal radiology dataset. Then, in collaboration with radiologists, we curate a VQA benchmark of clinically relevant multiple-choice problems designed to probe the model’s ability to distinguish between such images. By design, relying solely on unimodal (language) priors cannot achieve better than random guessing accuracy on our benchmark, and therefore performance on MediConfusion directly correlates with multimodal medical reasoning and image understanding capabilities. Remarkably, we discover that both state-of-the-art medical MLLMs, as well as the most advanced proprietary models, are easily confused by the image pairs, resulting in performance below random guessing for all models at the time of writing this paper. What is striking about this poor performance is that for some of the models (i.e. all medical MLLMs we studied) the images and corresponding captions are part of the training data! [^1] Given the public nature of the original dataset such images and captions are also likely part of the pre-training of proprietary models such as OpenAI’s GPT-4o, Google Deep Mind’s Gemini 1.5 Pro, and Anthropic’s Claude 3 Opus. Finally, we leverage our pipeline to categorize failure cases in order to guide future research toward more reliable medical AI solutions.
@@ 2 The MediConfusion Benchmark
The majority of existing multimodal foundation models leverage CLIP (Radford et al., 2021) to encode the input image (Li et al., 2024; Liu et al., 2024; Moor et al., 2023; Li et al., 2023). CLIP has been pretrained on internet-scale general domain data and, therefore, may not be suitable for the nuanced representation of medical images due to the considerable distribution shift. Thus, variants trained on large-scale medical image-text datasets have been introduced as image encoders for medical agents, including BiomedCLIP (Zhang et al., 2023b) and PMC-CLIP (Lin et al., 2023). Due to the specialized training data, these models are able to better capture the structure and semantics of medical images. However, surprisingly, we observe that the feature space of even specialized medical encoders is often not rich enough to clearly differentiate between images that are otherwise highly dissimilar. As a result, medical MLLM that leverage such input features suffer from impaired image understanding and visual reasoning, casting serious doubt on the reliability of such models in critical medical diagnosis. Therefore, designing challenging benchmarks that stress-test the visual capabilities of medical MLLM is of utmost importance for gaining a better understanding of the limitations of existing models.
@@ 2.1 Discovering confusing pairs
We find confusing image pairs in ROCO (Pelka et al., 2018), a multimodal dataset of $\approx 80k$ radiology images and their corresponding captions extracted from PMC-OA (Lin et al., 2023) (Figure 1, left). Inspired by Tong et al. (2024), we seek out pairs with clear visual differences, but high similarity in the feature space of the medical MLLMs. This implies that at least one of the images in the pair is compressed ambiguously, and thus, it is likely that relevant visual information is lost in the encoding. In particular, we base our selection criteria on BiomedCLIP’s embedding space, as this model has been specifically trained on medical image-text data, and thus it has a more refined feature space for medical images than a general-domain encoder, such as CLIP. Simply put, a radiology image pair that confuses BiomedCLIP, will likely confuse CLIP as well. Moreover, BiomedCLIP has been pretrained on the largest dataset of medical image-caption data among publicly available CLIP-style biomedical vision-language models. We measure visual differences between images in the feature space of DINOv2, a state-of-the-art vision-only foundation model with robust image representations that capture visual details. We randomly sample pairs of images and evaluate their similarity in BiomedCLIP ($sim_{med}$) and DINOv2 ($sim_{gen}$) feature spaces. We consider them a confusing pair if $sim_{med}\geq 0.9$ and $sim_{gen}\leq 0.75$ hold at the same time. The gap $|sim_{med}-sim_{gen}|$ can be increased further in order to obtain more difficult pairs, however we find that our setting is already challenging enough for most contemporary models. We depict sample pairs uncovered by our technique in Figure 2.
@@ 2.2 VQA generation
Figure 3: A sample question pair from MediConfusion. A confusing pair shares the same question and corresponding answer options, but the correct answer is different for the two (A for the image on the left and B for the image on the right). The model receives a set score of 1 only if it correctly answers both questions in the confusing pair. Individual score is evaluated separately for each image ($1$ out of $2$ in the example).
@@ 2.3 Data filtering and revision via radiologist feedback
As we generate the questions/answer choices using an LLM, various issues may arise such as factual errors and inconsistencies in quality, format, or language. To ensure the curation of a reliable benchmark dataset, oversight and feedback from a radiology expert are crucial. A radiologist evaluated each of the automatically generated VQA problems focusing on three aspects.
@@
Based on these guidelines, the radiologist assigned a quality score to each question, on a scale $1-10$. Higher scores correspond to better problems, and a score of $1$ is assigned if correctness is violated in any form (e.g., irrelevant question, incorrect answer). We add a VQA pair to MediConfusion only if the quality score is at least $5$ for both individual problems in the pair. Moreover, the expert verified the medical categories assigned by the LLM to each of the images and revised the question and answer options to improve language quality and precision. This step is crucial in eliminating model artifacts that originate in LLM-generated text inputs.
@@ 3.1 Evaluation
Prefix-based score (PS) – Following Xu et al. (2023), we compute the normalized likelihood of image-question-answer triplets for each answer option, and pick the option with the highest likelihood as the answer. In other words, we select the answer option that the model assigns the highest probability to, given the image and question. To compute the prefix-based score, we concatenate the medical question and the answer sentence directly, stripping the multiple choice question style (e.g., removing "Choose the letter of the correct answer.") and option indicators (e.g., removing "Option A: …") in order to ensure that models that have not been specifically trained on multiple choice question answering can also consistently provide valid answers.
We evaluate a representative set of $9$ models, $3$ of which are medical MLLMs (LLaVA-Med (Li et al., 2024), Med-Flamingo (Moor et al., 2023), RadFM (Wu et al., 2023b)), $3$ are flagship proprietary models (GPT-4o, Claude 3 Opus, Gemini 1.5 Pro) and $3$ open-source general-domain MLLMs (LLaVA-7B v1.6 (Liu et al., 2024), BLIP-2 (Li et al., 2023), InstructBLIP (Dai et al., 2023)). We set generation parameters according to the corresponding code release and recommended settings and use few-shot prompting for Med-Flamingo (more details in Appendix B).
3.2 Results
We summarize the performance of MLLMs on MediConfusion in Table 1. Alarmingly, all MLLMs perform below random guessing in terms of set accuracy, corroborating our hypothesis that models struggle to differentiate in fine enough detail between the extracted image pairs necessary for accurate medical reasoning. This observation is further supported by the markedly high (often above $90\%$) confusion scores indicating that models tend to select the same answer for both images within a confusing pair. Even RadFM, a model that does not leverage a CLIP-style image encoder, is confused on our benchmark ($82.39\%$ confusion score) with performance well below random guessing. As most likely proprietary models leverage visual encoders other than CLIP as well, the overall poor performance and extremely high confusion scores suggest that the exposed vulnerability is more general and not solely rooted in the specific ambiguities of CLIP encoding.
An interesting outlier is Gemini 1.5, which has been the least confused (approx. $60\%$) on the dataset; however, its accuracy is still close to random guessing. This may suggest that the model’s visual representations are rich enough to meaningfully distinguish between images; however, the medical knowledge or necessary reasoning skills are lacking to correctly answer the questions.
Furthermore, perhaps surprisingly, medical MLLMs did not outperform other methods, which indicates that the shortcomings cannot be addressed exclusively by domain-specific training. These results are especially surprising, as the image-caption pairs used to generate MediConfusion are part of PMC-OA, which is included in the pre-training set of all $3$ medical MLLMs in our experiments. We also note that given the public nature of PMC-OA these image-caption pairs are likely included in the training set of proprietary models as well. Finally, we do see some performance gap between open-source and proprietary models, with GPT-4o achieving the highest individual accuracy, however, barely surpassing that of random guessing ($56.25\%$ vs. $50\%$) with set accuracy still well below random guessing.
@@ 3.2 Results
We further break down the results based on the category of the question in order to identify if specific areas have been more/less challenging to the models. We summarize our findings in Table 2. Even though the overall results across all categories are close to random guessing performance, proprietary models demonstrate slightly better accuracies on questions related to cerebral and vascular images. In particular, GPT-4o achieves $34.25\%$ and $67.12\%$ set and individual accuracy correspondingly on vascular images, an overall best across all models and categories
@@
[…]
@@ 4.1 Identifying Patterns in Confusing Pairs
  • Pattern 1: Normal/variant anatomy vs. pathology– Models often struggle with differentiating between normal/variant anatomy and pathological structures. For instance, the model often confuses malalignment with normal alignment (e.g., atlantoaxial dislocation vs. normal atlantoaxial interval) or differentiating pituitary region masses (suprasellar vs. parasellar vs. intrasellar) or various anatomical regions of the spine (cervical vs. thoracic vs. lumbar).
@@
Most of the above shortcomings can be, to some degree, traced back to known, common failure modes (Tong et al., 2024) of visual reasoning in MLLMs in the general domain.
Detecting presence (or absence) of specific features: Correct reasoning over medical VQA problems strongly relies on detecting the presence (or absence) of particular features or objects relevant to the question. MLLMs are known to suffer from object hallucinations (Li et al., ) rooted in parts in flawed image encoding, statistical biases and strong reliance on language priors (Leng et al., 2024). We can see this specific weakness reflected in Patterns 3 and 4 directly.
Understanding state and condition: In medical VQA, it is crucially important for the model to understand the difference between "normal" and "abnormal" structures. MLLMs have difficulties identifying the state and condition of objects in the general domain, such as whether the ground is wet or if a flag is blowing in the air (Tong et al., 2024). These challenges may be amplified in the more nuanced medical setting, which we observe in Patterns 1 and 3 especially.
Positional and relational context: Answering medical VQA problems often necessitate a careful understanding of the spatial relationships of various anatomical features and their specific location. Recent research has uncovered serious limitations in the spatial reasoning capabilities of MLLMs (Kamath et al., 2023), some even failing to distinguish left from right. This pervasive weakness in spatial reasoning may translate to failures in medical VQA seen in Pattern 1.
Color and appearance: Recent work has shown that MLLMs can confuse colors and their intensity (bright/dark) (Tong et al., 2024), which may cause challenges in identifying signal characteristics in radiology images (high/low intensity) reflected in Pattern 2.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have tremendous potential to improve the accuracy, availability, and cost-effectiveness of healthcare by providing automated solutions or serving as aids to medical professionals. Despite promising first steps in developing medical MLLMs in the past few years, their capabilities and limitations are not well understood. Recently, many benchmark datasets have been proposed that test the general medical knowledge of such models across a variety of medical areas. However, the systematic failure modes and vulnerabilities of such models are severely underexplored with most medical benchmarks failing to expose the shortcomings of existing models in this safety-critical domain. In this paper, we introduce MediConfusion, a challenging medical Visual Question Answering (VQA) benchmark dataset, that probes the failure modes of medical MLLMs from a vision perspective. We reveal that state-of-the-art models are easily confused by image pairs that are otherwise visually dissimilar and clearly distinct for medical experts. Strikingly, all available models (open-source or proprietary) achieve performance below random guessing on MediConfusion, raising serious concerns about the reliability of existing medical MLLMs for healthcare deployment. We also extract common patterns of model failure that may help the design of a new generation of more trustworthy and reliable MLLMs in healthcare. The dataset and evaluation code are available at https://github.com/AIF4S/MediConfusion.
footnotetext: Published in International Conference on Learning Representations (ICLR) 2025.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have demonstrated unprecedented capabilities in a variety of multimodal tasks, including image understanding and visual reasoning, autonomous driving [6], robotics [38] and embodied AI [9]. Motivated by this success, a growing body of work [27, 18, 22] explores the potential of MLLMs in medical applications with the hope of paving the way to more accurate, personalized and cost-effective healthcare solutions through modern generative AI.
@@
Even though MLLMs show enormous potential in a wide range of tasks, a swath of challenges have stymied their deployment, including object hallucinations [21], relationship hallucinations [43], inaccurate object counting [14] and lack of spatial reasoning capabilities [15]. These shortcomings are especially worrisome in safety-critical applications, such as healthcare, where reliability is an essential requirement. In fact, recent research efforts on medical MLLMs have revealed weak anatomic knowledge [28], concerns on toxicity and patient privacy [45], highly unreliable disease diagnosis [41], and the fact that even a junior doctor far outperforms the most proficient medical MLLMs across a wide spectrum of tasks [39]. As model failure in the medical domain can lead to serious adverse health effects, it is of utmost importance to understand the performance and limitations of generative AI in the medical context.
@@
A flurry of activity has emerged around probing the performance of medical MLLMs in a multitude of tasks, including visual question answering (VQA) [2], disease classification and report generation [32], and modality recognition [42]. Even though the proposed medical benchmarks offer valuable insights on model performance across a variety of anatomic regions and imaging modalities, they are focused on evaluating the medical knowledge of MLLMs across large evaluation sets, heavily biased towards common or typical scenarios. Therefore, it is unclear how well the measured performance correlates with the actual multimodal medical reasoning capabilities of these models, especially in the face of systematic but perhaps more intricate model failures underrepresented in the dataset. Therefore, developing new evaluation benchmarks that carefully test and probe the capabilities of these systems, expose their vulnerabilities, and facilitate the development of a better understanding of failure modes is vital in healthcare applications.
@@
In this work, we introduce MediConfusion, a challenging benchmark for evaluating the failure modes of medical MLLMs from a vision perspective. We combine novel insights on the image representations of medical MLLMs with the expertise of clinical radiologists to craft a benchmark dataset designed to evaluate the ability of state-of-the-art models to recognize subtle, yet clinically meaningful differences between medical images. Our work reveals that medical MLLMs often confuse image pairs that otherwise appear very different in the image domain. Leveraging this observation, we introduce an automated pipeline to discover such pairs in the ROCO [30] multimodal radiology dataset. Then, in collaboration with radiologists, we curate a VQA benchmark of clinically relevant multiple-choice problems designed to probe the model’s ability to distinguish between such images. By design, relying solely on unimodal (language) priors cannot achieve better than random guessing accuracy on our benchmark, and therefore performance on MediConfusion directly correlates with multimodal medical reasoning and image understanding capabilities. Remarkably, we discover that both state-of-the-art medical MLLMs, as well as the most advanced proprietary models, are easily confused by the image pairs, resulting in performance below random guessing for most models at the time of writing this paper. What is striking about this poor performance is that for some of the models (i.e. all medical MLLMs we studied) the images and corresponding captions are part of the training data! [^1] Given the public nature of the original dataset such images and captions are also likely part of the pre-training of proprietary models such as OpenAI’s GPT-4o, Google Deep Mind’s Gemini 1.5 Pro, and Anthropic’s Claude 3 Opus. Finally, we leverage our pipeline to categorize failure cases in order to guide future research toward more reliable medical AI solutions.
@@ 2 The MediConfusion Benchmark
The majority of existing multimodal foundation models leverage CLIP [31] to encode the input image [18, 25, 27, 20]. CLIP has been pretrained on internet-scale general domain data and, therefore, may not be suitable for the nuanced representation of medical images due to the considerable distribution shift. Thus, variants trained on large-scale medical image-text datasets have been introduced as image encoders for medical agents, including BiomedCLIP [50] and PMC-CLIP [22]. Due to the specialized training data, these models are able to better capture the structure and semantics of medical images. However, surprisingly, we observe that the feature space of even specialized medical encoders is often not rich enough to clearly differentiate between images that are otherwise highly dissimilar. A growing body of recent work [34, 48] has shown that the contrastive pretraining objective, shared by common image encoders for MLLMs, can be optimized via shortcuts that lead to fundamental flaws in multimodal understanding. In particular, training consists of aligning image features with their corresponding text features within a batch of data. Thus, if the images are clearly distinct within the batch, the task becomes easy and the model is not encouraged to learn embeddings nuanced enough for more intricate downstream tasks, such as medical reasoning. As a result, MLLMs that leverage such pretrained encoders suffer from impaired image understanding and visual reasoning [35, 36], casting serious doubt on the reliability of such models in critical medical diagnosis. Therefore, designing challenging benchmarks that stress-test the visual capabilities of medical MLLM is of utmost importance for gaining a better understanding of the limitations of existing models.
@@ 2.1 Discovering confusing pairs
We find confusing image pairs in ROCO [30], a multimodal dataset of $\approx 80k$ radiology images and their corresponding captions extracted from PMC-OA [22] (Figure 1, left). Inspired by Tong et al. [36], we seek out pairs with clear visual differences, but high similarity in the feature space of medical vision-language models. This implies that at least one of the images in the pair is compressed ambiguously, and thus, it is likely that relevant visual information is lost in the encoding. In particular, we base our selection criteria on BiomedCLIP’s embedding space, as this model has been specifically trained on medical image-text data, and thus it has a more refined feature space for medical images than a general-domain encoder, such as CLIP. Simply put, a radiology image pair that confuses BiomedCLIP, will likely confuse CLIP as well. Moreover, BiomedCLIP has been pretrained on the largest dataset of medical image-caption data among publicly available CLIP-style biomedical vision-language models. We measure visual differences between images in the feature space of DINOv2, a vision-only foundation model with robust image representations that capture visual details. We randomly sample pairs of images and evaluate their similarity in BiomedCLIP ($sim_{med}$) and DINOv2 ($sim_{gen}$) feature spaces. We consider them a confusing pair if $sim_{med}\geq 0.9$ and $sim_{gen}\leq 0.75$ hold at the same time. The gap $|sim_{med}-sim_{gen}|$ can be increased further in order to obtain more difficult pairs, however we find that our setting is already challenging enough for most contemporary models. We depict sample pairs uncovered by our technique in Figure 2.
@@ 2.2 VQA generation
Figure 3: A VQA pair from MediConfusion. A confusing pair shares the same question and answer options, but the correct answer is different for the two (A for the image on the left and B for the image on the right). The model receives a set score only if it correctly answers both questions in the confusing pair. Individual score is evaluated separately for each image.
@@ 2.3 Data filtering and revision via radiologist feedback
As we generate the questions/answer choices using an LLM, various issues may arise such as factual errors and inconsistencies in quality, format, or language. To ensure the curation of a reliable benchmark dataset, oversight and feedback from a radiology expert are crucial. A radiologist has evaluated each of the automatically generated VQA problems focusing on three aspects.
@@
Based on these guidelines, the radiologist has assigned a quality score to each question, on a scale $1-10$. Higher scores correspond to better problems, and a score of $1$ is assigned if correctness is violated in any form (e.g., irrelevant question, incorrect answer). We add a VQA pair to MediConfusion only if the quality score is at least $5$ for both individual problems in the pair. Moreover, the expert has verified the medical categories assigned by the LLM to each of the images and revised the question and answer options to improve language quality and precision. This step is crucial in eliminating model artifacts that originate in LLM-generated text inputs.
@@ 3.1 Evaluation
Prefix-based score (PS) – Following Xu et al. [46], we compute the normalized likelihood of image-question-answer triplets for each answer option, and pick the option with the highest likelihood as the answer. In other words, we select the answer option that the model assigns the highest probability to, given the image and question. To compute the prefix-based score, we concatenate the medical question and the answer sentence directly, stripping the multiple choice question style (e.g., removing "Choose the letter of the correct answer.") and option indicators (e.g., removing "Option A: …") in order to ensure that models that have not been specifically trained on multiple choice question answering can also consistently provide valid answers.
We evaluate a representative set of $13$ models, $3$ of which are medical MLLMs (LLaVA-Med [18], Med-Flamingo [27], RadFM [42]), $5$ are flagship proprietary models (GPT-4o, o1, Claude 3 Opus, Gemini 1.5 Pro, Gemini 2.0 Flash) and $5$ open-source general-domain MLLMs (LLaVA-7B v1.6 [25], BLIP-2 [20], InstructBLIP [7], DeepSeek-VL2[44], Molmo-7B [8]). We set generation parameters according to the corresponding code release and recommended settings and use few-shot prompting for Med-Flamingo (more details in Appendix B).
[…]
@@
Gemini 2.0 Flash 29.11 64.56 39.73 68.49 19.40 58.21 23.53 58.86 40.48 66.67 26.92 61.54 25.58 60.47 30.00 50.00 42.86 71.43
3.2 Results
We summarize the performance of MLLMs on MediConfusion in Table 1. Alarmingly, almost all MLLMs perform below random guessing in terms of set accuracy, corroborating our hypothesis that models struggle to differentiate in fine enough detail between the extracted image pairs necessary for accurate medical reasoning. This observation is further supported by the markedly high (often above $90\%$) confusion scores, indicating that models tend to select the same answer for both images within a confusing pair. Even RadFM, a model that does not leverage a CLIP-style image encoder, is confused on our benchmark ($82.39\%$ confusion score) with performance well below random guessing. As most likely proprietary models leverage visual encoders other than CLIP as well, the overall poor performance and extremely high confusion scores suggest that the exposed vulnerability is more general and not solely rooted in the specific ambiguities of CLIP-style contrastive pretraining.
Gemini models are interesting outliers: even though they are the least confused on the dataset, they struggle tackling MediConfusion as well, as Gemini 2.0 barely surpasses random guessing by about $3\%$. This may suggest that the model’s visual representations are rich enough to meaningfully distinguish between images; however, the medical knowledge or necessary reasoning skills are lacking to correctly answer the questions.
Furthermore, perhaps surprisingly, medical MLLMs did not outperform other methods, indicating that the shortcomings cannot be addressed exclusively by domain-specific training. These results are especially surprising, as the image-caption pairs used to generate MediConfusion are part of PMC-OA, which is included in the pre-training set of all medical MLLMs in our experiments. We also note that given the public nature of PMC-OA these image-caption pairs are likely included in the training set of proprietary models as well. Finally, we see some performance gap between open-source and proprietary models, with Gemini $2.0$ achieving the highest individual accuracy of $61.93\%$.
We further break down the results based on the category of the question in order to identify if specific areas have been more/less challenging to the models. We summarize our findings in Table 2. Even though the overall results across all categories are close to random guessing performance, proprietary models demonstrate slightly better accuracies on questions related to cerebral, vascular, and nuclear medicine images. In particular, Gemini $2.0$ achieves $42.86\%$ and $71.43\%$ set and individual accuracy correspondingly on nuclear medicine, an overall best across all models and categories
@@ 4.1 Identifying Patterns in Confusing Pairs
  • Pattern 1: Normal/variant anatomy vs. pathology– Models often struggle to differentiate between normal/variant anatomy and pathological structures. For instance, the model often confuses malalignment with normal alignment (e.g., atlantoaxial dislocation vs. normal atlantoaxial interval) or differentiating pituitary region masses (suprasellar vs. parasellar vs. intrasellar) or various anatomical regions of the spine (cervical vs. thoracic vs. lumbar).
@@
Most of the above shortcomings can be, to some degree, traced back to known, common failure modes [36] of visual reasoning in MLLMs in the general domain.
Detecting presence (or absence) of specific features: Correct reasoning over medical VQA problems strongly relies on detecting the presence (or absence) of particular features or objects relevant to the question. MLLMs are known to suffer from object hallucinations [21] rooted in parts in flawed image encoding, statistical biases and strong reliance on language priors [17]. We can see this specific weakness reflected in Patterns 3 and 4 directly.
Understanding state and condition: In medical VQA, it is crucially important for the model to understand the difference between "normal" and "abnormal" structures. MLLMs have difficulties identifying the state and condition of objects in the general domain, such as whether the ground is wet or if a flag is blowing in the air [36]. These challenges may be amplified in the more nuanced medical setting, which we observe in Patterns 1 and 3 especially.
Positional and relational context: Answering medical VQA problems often necessitate a careful understanding of the spatial relationships of various anatomical features and their specific location. Recent research has uncovered serious limitations in the spatial reasoning capabilities of MLLMs [15], some even failing to distinguish left from right. This pervasive weakness in spatial reasoning may translate to failures in medical VQA seen in Pattern 1.
Color and appearance: Recent work has shown that MLLMs can confuse colors and their intensity (bright/dark) [36], which may cause challenges in identifying signal characteristics in radiology images (high/low intensity) reflected in Pattern 2.